OpenStudy (anonymous):

By what number must you multiply each side of the equation 2/3x=10 to produce the equivalent equation x=15 ? A.3/2 B.-2/3 C.3 D.1/3

OpenStudy (anonymous):

@izuru

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@madhu.mukherjee.946

OpenStudy (anonymous):

can you help me @Vocaloid ?

OpenStudy (johnweldon1993):

Hint \[\large \frac{a}{bx} = c \rightarrow x= \frac{b}{a}(c)\]

OpenStudy (anonymous):

and then use 10 or 15 to multiply the 2 numbers?

OpenStudy (johnweldon1993):

We are just asked to turn \(\large \frac{2}{3x} = 10\) into \(\large x = 15\) So we needed to isolate 'x' which can be done by...? Look at my above post

OpenStudy (anonymous):

b/a 15 or b/a 10?

OpenStudy (johnweldon1993):

Compare the two \[\large \frac{2}{3x} = 10 \rightarrow \frac{a}{bx} = c\] \[\large x = ?\]

OpenStudy (anonymous):

x=15

OpenStudy (johnweldon1993):

Yes we know that 'x = 15' that is given Look again what I posted, I'll copy it all at once \[\large \frac{a}{bx}=c \rightarrow x = \frac{b}{a}(c)\] You have \[\large \frac{2}{3x} = 10 \rightarrow x = \frac{?}{?}(10)\]

OpenStudy (madhu.mukherjee.946):

i think its option A coz if you multiply 3/2 on both sides you get x=15